### v1

# Enhancing Clinical Trial Patient Matching through Knowledge Augmentation with Multi-Agents

Abstract
Matching patients effectively and efficiently for clinical trials is a significant challenge due to the complexity and variability of patient profiles and trial criteria. This paper presents a novel framework, Multi-Agents for Knowledge Augmentation (MAKA), designed to enhance patient-trial matching by dynamically supplementing matching prompts with external, domain-specific knowledge. The MAKA architecture consists of five key components: a knowledge probing agent that detects gaps in domain knowledge, a navigation agent that manages interactions among multiple specialized knowledge augmentation agents, a knowledge augmentation agent that incorporates relevant information into patient-trial matching prompts, a supervision agent aligning the outputs from other agents with the instructions and a matching agent making the final selection decision. This approach enhances the accuracy and contextual richness of patient matching, addresses inherent knowledge gaps in both trail criteria and large language models (LLMs), and improves the alignment between patient characteristics and the criteria.
1 Introduction
Patient matching for clinical trials refers to the process of identifying and enrolling participants whose health profiles align with the specific eligibility criteria of a given clinical study. Efficient patient matching is crucial to accelerating the drug development process and reducing the overall cost of clinical trials. It provides an opportunity for participants to access experimental treatments that could potentially improve their health outcomes and, in some cases, transform their quality of life.
Patient matching remains a challenging task despite its critical role in clinical trials. Traditional automated methods, which are mostly rule-based, depend heavily on hand-crafted rules and are often time-consuming (Brøgger-Mikkelsen et al., 2020). These approaches are also highly sensitive to variations in how patient and trial data are expressed in natural language across different sources. Simple machine learning models often struggle with the load of information and the low density of relevant sentences in lengthy electronic health record(EHR) data. They usually underperform when dealing with complex eligibility criteria or inconsistent data formats, resulting in inefficiencies and mismatches that can negatively impact trial outcomes (Kadam et al., 2016).
The emergence of of Large Language Models (LLMs) offers a promising opportunity to improve patient matching. LLMs, such as GPT-4 Achiam et al. (2023), usually possess a large text window and advanced natural language understanding capabilities (Karanikolas et al., 2023), which enable them process complex trail criteria and patient data effectively. However, their general knowledge may not always be sufficient for knowledge-intensive tasks like patient matching in clinical trials, where domain-specific expertise is required. The knowledge gap inherent in LLMs and misssing information in the pre-defined criteria can limit their performance, highlighting the need for LLM augmentation of specialized medical knowledge (Shuster et al., 2021). While research indicates that LLMs do encode medical knowledge in a parameterized form (Singhal et al., 2023), this information can be outdated or misaligned with the special needs of certain clinical trials.
To address these limitations, we present a new framework called: Multi-Agents for Knowledge Augmentation (MAKA). MAKA is a novel architecture designed to enhance the patient matching process by dynamically integrating domain-specific knowledge aligned with the unique context of clinical trials. It consists of five main components: the Knowledge Probing Agent, Navigation Agent, Knowledge Augmentation Agent, Supervision Agent and Matching Agent. By leveraging this multi-agent framework, MAKA effectively identifies and fills knowledge gaps, enhancing the LLM’s ability to understand complex eligibility criteria and patient profiles. This leads to an average improvement of 3% on accuracy and 6% on F1-score across various patient selection criteria, with up to a 10% boost in more knowledge-intensive cases. Overall, our framework shows significant potentials for achieving more accurate patient matching.
2 Related Work
2.1 Clinical Trail Patient Matching
Traditional patient matching systems rely on query-based structured databases, such as Criteria2Query Yuan et al. (2019), or handcrafted rule-based approaches like RBC Oleynik et al. (2019). However, these methods struggle to handle unstructured free-text data and are highly sensitive to variations in formatting and verbal expressions. Unfortunately, a significant portion of EHRs consists of free-text, including clinical notes, surgical records, radiology reports, pathology reports and so on Sun et al. (2018). Thus, the effectiveness and scalability of traditional methods are significantly limited, making them unsuitable for future patient matching tasks.
Recent advancements in LLMs have shown remarkable capabilities in directly processing free-text clinical trial criteria, offering scalable solutions without the need for extensive rule-based preprocessing. Many Studies have applied these models to patient-centric trial matching tasks using real-world data. For instance, Nievas et al. (2024) fine-tuned the open-source Llama 2 Touvron et al. (2023) to automate patient matching. However, fine-tuning LLMs is computationally expensive, prompting alternative approaches that leverage frozen models. Jin et al. (2023) introduced TrialGPT, a patient-centric matching framework that assesses patient eligibility on a criterion-by-criterion basis, filtering out ineligible trials and ranking potentially eligible ones. Their results indicated a strong correlation between TrialGPT’s predictions and expert clinician judgments, achieving accuracy at the criterion level similar to human experts. Similarly, Gupta et al. (2024) proposed OncoLLM, an end-to-end pipeline that converts trial criteria into conversational prompts and uses a Chain-of-Thought (CoT) Wei et al. (2022) reasoning strategy. OncoLLM outperforms GPT-3.5 while underperforms GPT-4, despite its smaller model size.
In contrast, other studies have focused on trial-centric patient matching, typically using public deidentified datasets due to HIPAA-related data constraints. Wornow et al. (2024) developed a LLM-based system with a two-stage retrieval pipeline to identify suitable patients for trials using a zero-shot prompt approach. It reduces token processing costs while outperforming existing state-of-the-art models on the n2c2 2018 cohort selection benchmark Stubbs et al. (2019). Likewise, Beattie et al. (2024) applied a CoT strategy in a Retrieval-Augmented Generation (RAG) Lewis et al. (2020) system using GPT-3.5 Turbo and GPT-4, demonstrating results comparable to those of Wornow et al. (2024).
These works collectively highlight the potential of LLMs in clinical trial matching. However, existing approaches directly employ LLMs without verifying whether the trial criteria are sufficiently understood by the model Webson and Pavlick (2021) or whether there are gaps in the model’s knowledge Lewis et al. (2020). Our work fits within the trial-centric patient matching literature and makes a novel contribution by automatically identifying and addressing these gaps, thereby enhancing the accuracy patient matching.
2.2 LLM-based Multi-Agents (LLM-MA)
LLM-MA have seen rapid development across various domains, driven by the promising potential of LLMs as autonomous agents. Guo et al. (2024) provide a comprehensive overview of recent advancements in LLM-MA applications, categorizing them into two primary areas: Problem-Solving and World Simulation. Problem-solving LLM-MA systems explore the capabilities of agents with specialized expertise, enabling them to outperform single-agent LLM systems through collaborative efforts. For instance, Wang et al. (2024) proposed Mixture-of-Agents, a layered multi-agent architecture that achieves state-of-the-art performance in benchmarks like AlpacaEval 2.0 Dubois et al. (2024). Similarly, Hong et al. (2023) introduced MetaGPT, a meta-programming framework that integrates human-like task decomposition and workflows, outperforming GPT-4 in various code generation tasks.
In the healthcare domain, the use of LLM-MA systems is emerging. TRIAGEAGENT Lu et al. (2024) utilizes LLMs for role-playing interactions, incorporating self-confidence and early-stopping mechanisms in multi-round discussions to improve document reasoning and classification precision for triage tasks. Tang et al. (2023) uses multiple LLM-based agents as domain experts to collaboratively analyze medical reports and reach consensus on medical diagnoses. To the best of our knowledge, this is the first study to leverage LLM-MA for cohort selection in clinical trials. Our framework leverages LLM-MA agency to dynamically augment patient-trial matching with tailored knowledge, thereby enriching the contextual understanding of patient profiles and achieving better patient matching results.
3 Methods
We automate decomposed patient-matching workflows by utilizing the ability of LLMs to process and generate complex natural language. The proposed framework, illustrated in Figure 1, comprises five interconnected components, each with a distinct role in ensuring robust automation.
Knowledge Probing Agent: This component identifies knowledge gaps, both in terms of missing information in criteria and deficiencies in the LLM’s understanding of clinical trials and patient characteristics. By employing prompt-based self-probing techniques, inspired by SELFREF Kadavath et al. (2022) and MOREINFO Feng et al. (2023), the agent evaluates its own responses for correctness and consistency.If the agent determines that the original criterion is detailed and understandable enough for other agents, the criterion will be directly forwarded to the Matching Agent. Otherwise, the criterion is forwarded to the Navigation Agent for further processing.
Navigation Agent: In the MAKA framework, multiple specialized augmentation agents are available to enrich the criterion with domain-specific knowledge. The Navigation Agent determines the most suitable augmentation agent for the task and routes the workflow accordingly.
Knowledge Augmentation Agent: This component enriches the criterion with relevant, domain-specific information from different source. For instance, the Retrieval Agent Lewis et al. (2020) retrieves information from indexed medical databases, the Online Search Agent gathers supplementary knowledge from online sources, and the Self-Augment Agent leverages the LLM’s intrinsic knowledge to refine the original criterion.
Supervision Agent: Acting as a supervisor, this agent ensures that the augmented criterion aligns with the original instructions. For example, the augmented output must retain verbatim elements from the original criterion. This supervisory mechanism enhances the framework’s instruction-following capability and maintains consistency.
Matching Agent: Finally, with either the augmented or unaugmented criterion, as determined by the Knowledge Probing Agent, the Matching Agent evaluates patient eligibility in a zero-shot manner.
4 Evaluation
| Criteria | Accuracy | Precision | Recall | F1-score | ||||||||
| Zero-shot | CoT | MAKA | Zero-shot | CoT | MAKA | Zero-shot | CoT | MAKA | Zero-shot | CoT | MAKA | |
| ABDOMINAL | 0.858 | 0.868 | 0.847 | 0.893 | 0.897 | 0.944 | 0.701 | 0.729 | 0.626 | 0.785 | 0.804 | 0.753 |
| ADVANCED-CAD | 0.854 | 0.854 | 0.868 | 0.895 | 0.872 | 0.929 | 0.853 | 0.882 | 0.841 | 0.873 | 0.877 | 0.883 |
| ALCOHOL-ABUSE | 0.969 | 0.969 | 0.976 | 0.529 | 0.545 | 0.667 | 0.900 | 0.600 | 0.600 | 0.667 | 0.571 | 0.632 |
| ASP-FOR-MI | 0.861 | 0.875 | 0.885 | 0.899 | 0.908 | 0.919 | 0.930 | 0.939 | 0.939 | 0.915 | 0.923 | 0.929 |
| CREATININE | 0.802 | 0.802 | 0.885 | 0.664 | 0.667 | 0.812 | 0.934 | 0.925 | 0.896 | 0.776 | 0.775 | 0.852 |
| DIETSUPP-2MOS | 0.764 | 0.781 | 0.785 | 0.704 | 0.717 | 0.721 | 0.940 | 0.953 | 0.953 | 0.805 | 0.818 | 0.821 |
| DRUG-ABUSE | 0.941 | 0.920 | 0.993 | 0.469 | 0.395 | 0.882 | 1.000 | 1.000 | 1.000 | 0.638 | 0.566 | 0.938 |
| ENGLISH | 0.976 | 0.986 | 0.983 | 0.985 | 0.989 | 0.996 | 0.989 | 0.996 | 0.985 | 0.987 | 0.992 | 0.991 |
| HBA1C | 0.840 | 0.830 | 0.878 | 0.722 | 0.702 | 0.807 | 0.892 | 0.902 | 0.863 | 0.798 | 0.790 | 0.834 |
| KETO-1YR | 0.993 | 0.990 | 0.993 | - | - | - | - | - | - | - | - | - |
| MAJOR-DIABETES | 0.833 | 0.826 | 0.847 | 0.851 | 0.831 | 0.873 | 0.840 | 0.853 | 0.840 | 0.845 | 0.842 | 0.856 |
| MAKES-DECISIONS | 0.920 | 0.924 | 0.931 | 0.985 | 0.985 | 0.996 | 0.931 | 0.935 | 0.931 | 0.957 | 0.959 | 0.963 |
| MI-6MOS | 0.885 | 0.872 | 0.927 | 0.434 | 0.396 | 0.571 | 0.885 | 0.808 | 0.769 | 0.582 | 0.532 | 0.656 |
| Average | 0.884 | 0.883 | 0.909 | 0.727 | 0.715 | 0.822 | 0.894 | 0.870 | 0.846 | 0.785 | 0.769 | 0.828 |
4.1 Data Description
We use the dataset from Track 1 of the 2018 n2c2 challenge Stubbs et al. (2019), which is commonly used for evaluating patient matching in clinical trials. The dataset consists of 288 de-identified patients, each having between 2 to 5 unstructured longitudinal clinical notes in American English. The entire corpus contains 781,006 tokens, with an average of 2,711 tokens per patient. The challenge simulates a synthetic clinical trial with 13 predefined inclusion criteria, resulting in 3744 patient-criteria pairs in total. The detailed criteria can be found in Table 4. For each patient-criteria pair, a binary label is provided, indicating whether the patient "MET" or "NOT MET" the criterion. Two annotators with medical background labeled these data. The disagreements between the two annotators were resolved by an independent MD to establish the final ground truth labels.
4.2 Baseline and Metrics
Previous research has demonstrated the superiority of LLM-based methods over traditional rule-based systems; hence, this study focuses exclusively on LLM-based benchmarks. We include two approaches: the zero-shot method with hand-crafted criteria Wornow et al. (2024), as shown in Table 5, and a Chain-of-Thought (CoT) approach Beattie et al. (2024) using the original criteria. In line with prior literature, we report accuracy, precision, recall, and F1-score for this binary classification task.
Previous studies have primarily focused on evaluating classification performance at the individual criterion level, often overlooking the overall patient-trial assessment. In this study, we address this gap by examining patient-trial selection performance. Specifically, a patient is eligible for the trial only if all inclusion criteria are met. However, no patient in the n2c2 dataset satisfied all 13 criteria. To mitigate this, we synthesize a trail with selected criteria that had at least 100 patients meeting them, resulting in 28 eligible patients from a total cohort of 288. We then evaluated our method’s on this new synthetic trail in identifying eligible patients and reported the findings.
4.3 Results
4.3.1 Criterion-Level Performance
Table 1 shows the performance of three approaches—Zero-shot,CoT, MAKA—across multiple individual criteria. Our results indicate that while the CoT method slightly outperforms Zero-shot in Accuracy and Recall, MAKA consistently achieves higher scores, particularly in Accuracy (0.909) and Precision(0.822) on average. The F1-score of MAKA (0.828) suggests a balanced performance across criteria, although it falls slightly behind in a few specific cases such as "ABDOMINAL" and "English".
The MAKA-augmented criteria, presented in Table 6, demonstrate significant improvements over the baseline human-redefined criteria outlined in Table 5 from Wornow et al. (2024). These enhancements primarily focus on providing explicit clarifications and detailed expansions of key concepts, enabling LLMs to better align with trial requirements. For example, MAKA supplements normal ranges for criteria like "HBA1C" and "CREATININE" across diverse demographic groups. And they are also the criteria that benefit the most from MAKA’s augmentation. Additionally, MAKA provides a more comprehensive definition for criteria like "MAJOR-DIABETES," and use more widely recognized terminology, "heart attack", to explain clinical terms "Myocardial Infarction (MI)". These refinements enhance the ability of LLMs to match patients from clinical notes written in a variety of styles, resulting in notable performance gains for these criteria.
| Metric | Zero-shot | CoT | MAKA |
| Accuracy | 0.8924 | 0.9201 | 0.9306 |
| Precision | 0.4615 | 0.5676 | 0.6333 |
| Recall | 0.6429 | 0.7500 | 0.6786 |
| F1-Score | 0.5373 | 0.6462 | 0.6552 |
4.3.2 Trial-Level Performance
The result of trail-levle analysis can be found in Table 2. At the trial level, where a patient’s eligibility is determined by satisfying all selected criteria, MAKA outperforms both Zero-shot and CoT. MAKA achieves the highest Accuracy (0.9306), Precision (0.6333)and F1-score (0.6552) but underperform CoT on Recall. The trail-level results further validate MAKA’s capability in maintaining a higher balance between identifying eligible patients and minimizing patient-filtering load
Our findings suggest that MAKA effectively can effectively identifies and addresses knowledge gaps in LLMs for trial-centric patient matching. By leveraging agent-augmented criteria, MAKA enhances LLMs’ understanding of clinical trial requirements. And it achieves a better performance on both minimizing manual selection efforts and accurately identifying eligible patients.
5 Limitations
One significant limitation of this study lies in the dataset used for evaluation. The n2c2 dataset, the most commonly used, publicly available dataset for the cohort selection task, is relatively small in size and contains criteria that are neither particularly complex nor challenging for LLM-based processing. As a result, while MAKA is designed as a general framework capable of handling both simple and complex criteria, the dataset’s limitations restrict our ability to thoroughly evaluate its generalizability across a broader range of clinical trial requirements.
An additional observation from our experiments is that the Knowledge Probing Agent predominantly identified gaps in the criteria themselves—primarily issues of clarity or insufficient detail—rather than uncovering knowledge gaps in LLM. Moreover, the Navigation Agent consistently determined that all criteria could be sufficiently augmented by the Self-Augmentation Agent, with no need to employ external retrieval or online search augmentation. These findings suggest that the dataset’s simplicity may have limited the full utilization of MAKA’s modular capabilities.
6 Conclusion
In this paper, we introduced MAKA, a framework designed to automate decomposed patient-matching workflows by leveraging the advanced natural language understanding capabilities of LLMs. The framework’s modular design incorporates five key components—Knowledge Probing Agent, Navigation Agent, Knowledge Augmentation Agent, Supervision Agent, and Matching Agent—each contributing to systematic knowledge augmentation and robust decision-making in patient-trail matching settings.
We demonstrated that MAKA significantly improves patient-matching performance at both the criterion and trial levels. Compared to baseline approaches like Zero-shot and Chain of Thought (CoT), MAKA achieves notable gains in Accuracy, Precision, and F1-score, particularly when applied to complex or ambiguous criteria. By enriching clinical criteria with domain-specific insights, MAKA consistently outperforms alternatives while maintaining a balance between precision and recall. Further investigation of agent behaviors indicates that the performance improvements primarily stem from filling the knowledge gaps existing in the criteria, rather than in LLMs.
Our results highlight the practical benefits of augmenting LLM capabilities with a structured muti-agent workflow for knowledge enhancement. For example, MAKA-augmented criteria improves understanding of clinical trial requirements by aligning better with diverse patient note styles and heterogeneity of criterion. The results validate MAKA’s potential for real-world deployment in clinical trial workflows.
It is worth noting that the design of MAKA supports the future development of a human-AI collaboration system. For the trial-centric patient matching task, the volume of criterion-supervision performed by the supervisory agent is relatively small and can be efficiently handled by human experts with minimal time investment. Exploring how humans can further enhance and facilitate this system presents an interesting direction for future research.
To address the limitations of this study and validate MAKA’s generalizability, future work could also focus on evaluating the framework using more challenging datasets, such as those involving oncology trial matching which contains highly complex criteria or those with a cohort of larger size. Testing MAKA in thess contexts would not only provide a more rigorous assessment of its capabilities but also demonstrate its potential to address critical real-world challenges in clinical trial recruitment.
Acknowledgments
We gratefully acknowledge support from the OpenAI Research Access Program, which provided access to advanced language models that were instrumental in developing and evaluating the framework presented in this work. This support significantly facilitated the experiments and analyses conducted for this research.
Appendix A Criteria
A.1 N2C2 Criteria Statistics
| Criterion | Met Count | Not Met Count |
| ABDOMINAL | 107 | 181 |
| ADVANCED-CAD | 170 | 118 |
| ALCOHOL-ABUSE | 10 | 278 |
| ASP-FOR-MI | 230 | 58 |
| CREATININE | 106 | 182 |
| DIETSUPP-2MOS | 149 | 139 |
| DRUG-ABUSE | 15 | 273 |
| ENGLISH | 265 | 23 |
| HBA1C | 102 | 186 |
| KETO-1YR | 1 | 287 |
| MAJOR-DIABETES | 156 | 132 |
| MAKES-DECISIONS | 277 | 11 |
| MI-6MOS | 26 | 262 |
A.2 N2C2 Original Criteria
| Criteria | Definition |
| DRUG-ABUSE | Drug abuse, current or past. |
| ALCOHOL-ABUSE | Current alcohol use over weekly recommended limits. |
| ENGLISH | Patient must speak English. |
| MAKES-DECISIONS | Patient must make their own medical decisions. |
| ABDOMINAL | History of intra-abdominal surgery, small or large intestine resection, or small bowel obstruction. |
| MAJOR-DIABETES | Major diabetes-related complications, such as: amputation, kidney damage, skin conditions, retinopathy, nephropathy, neuropathy. |
| ADVANCED-CAD | Advanced cardiovascular disease, defined as two or more of the following: medications for CAD, history of myocardial infarction, angina, ischemia. |
| MI-6MOS | Myocardial infarction in the past 6 months. |
| KETO-1YR | Diagnosis of ketoacidosis in the past year. |
| DIETSUPP-2MOS | Taken a dietary supplement (excluding Vitamin D) in the past 2 months. |
| ASP-FOR-MI | Use of aspirin to prevent myocardial infarction. |
| HBA1C | Any HbA1c value between 6.5 and 9.5%. |
| CREATININE | Serum creatinine > upper limit of normal. |
A.3 Redefined Criteria
| Criteria | Definition |
| ABDOMINAL | History of intra-abdominal surgery. This could include any form of intra-abdominal surgery, including but not limited to small/large intestine resection or small bowel obstruction. |
| ADVANCED-CAD | Advanced cardiovascular disease (CAD). Defined as having 2 or more of the following: (a) Taking 2 or more medications to treat CAD (b) History of myocardial infarction (MI) (c) Currently experiencing angina (d) Ischemia, past or present. The patient must have at least 2 of these categories (a,b,c,d) to meet this criterion, otherwise the patient does not meet this criterion. For ADVANCED-CAD, be strict in your evaluation of the patient – if they just have cardiovascular disease, then they do not meet this criterion. |
| ALCOHOL-ABUSE | Current alcohol use over weekly recommended limits. |
| ASP-FOR-MI | Use of aspirin for preventing myocardial infarction (MI). |
| CREATININE | Serum creatinine level above the upper normal limit. |
| DIETSUPP-2MOS | Consumption of a dietary supplement (excluding vitamin D) in the past 2 months. To assess this criterion, go through the list of medications and supplements taken from the note. If a substance could potentially be used as a dietary supplement (i.e., it is commonly used as a dietary supplement, even if it is not explicitly stated as being used as a dietary supplement), then the patient meets this criterion. Be lenient and broad in what is considered a dietary supplement. For example, a ’multivitamin’ and ’calcium carbonate’ should always be considered a dietary supplement if they are included in this list. |
| DRUG-ABUSE | Current or past history of drug abuse. |
| ENGLISH | Patient speaks English. Assume that the patient speaks English, unless otherwise explicitly noted. If the patient’s language is not mentioned in the note, then assume they speak English and thus meet this criteria. |
| HBA1C | Any hemoglobin A1c (HbA1c) value between 6.5% and 9.5%. |
| KETO-1YR | Diagnosis of ketoacidosis within the past year. |
| MAJOR-DIABETES | Major diabetes-related complication. Examples of “major complication” (as opposed to “minor complication”) include, but are not limited to, any of the following that are a result of (or strongly correlated with) uncontrolled diabetes: • Amputation • Kidney damage • Skin conditions • Retinopathy • Nephropathy • Neuropathy. Additionally, if multiple conditions together imply a severe case of diabetes, then count that as a major complication. |
| MAKES-DECISIONS | Patient must make their own medical decisions. Assume that the patient makes their own medical decisions, unless otherwise explicitly noted. There is no information provided about the patient’s ability to make their own medical decisions, then assume they do make their own decisions and therefore meet this criteria. |
| MI-6MOS | Myocardial infarction (MI) within the past 6 months. |
A.4 MAKA-Augemented Criteria
| Criteria | Definition |
| ABDOMINAL | Criteria: History of intra-abdominal surgery, including small/large intestine resection or small bowel obstruction. |
| Explanation: Intra-abdominal surgery involves operations on organs within the abdominal cavity (e.g., stomach, intestines). Small bowel obstruction refers to blockages in the small intestine that may require surgery. | |
| ADVANCED-CAD | Criteria: Advanced cardiovascular disease (CAD) defined by having two or more of the following: (a) Taking 2 or more medications for CAD, (b) History of myocardial infarction (MI), (c) Current angina, or (d) Ischemia (past or present). |
| Explanation: CAD affects heart function due to blocked arteries. Definitions include: MI (heart attack), angina (chest pain), ischemia (reduced blood flow), and specific medications like statins or beta-blockers. A diagnosis of CAD alone does not qualify as advanced. | |
| ALCOHOL-ABUSE | Criteria: Current alcohol use exceeding weekly recommended limits. |
| Explanation: Weekly limits: up to 14 standard drinks. A standard drink is 14g of pure alcohol (12oz beer, 5oz wine, or 1.5oz spirits). | |
| ASP-FOR-MI | Criteria: Use of aspirin to prevent myocardial infarction (MI). |
| Explanation: Aspirin reduces MI risk by preventing clots but requires medical advice to balance benefits (risk reduction) and potential side effects (e.g., bleeding). | |
| CREATININE | Criteria: Serum creatinine level above the upper normal limit. |
| Explanation: Creatinine reflects kidney function. Normal levels vary by age and gender (e.g., 0.6-1.2 mg/dL for adult males). Elevated levels suggest potential kidney impairment. | |
| DIETSUPP-2MOS | Criteria: Consumption of a dietary supplement (excluding Vitamin D) in the past 2 months. |
| Explanation: Supplements include multivitamins, calcium carbonate, fish oil, or herbal supplements like ginseng. Excludes Vitamin D explicitly. | |
| DRUG-ABUSE | Criteria: Current or past drug abuse history. |
| Explanation: Drug abuse includes misuse of psychoactive substances, with potential consequences like health issues or legal problems. | |
| ENGLISH | Criteria: Patient must speak English. |
| Explanation: English proficiency ensures effective participation in English-based trials. Assume English proficiency unless noted otherwise. | |
| HBA1C | Criteria: Hemoglobin A1c (HbA1c) between 6.5% and 9.5%. |
| Explanation: HbA1c measures average blood sugar over 2–3 months. The range identifies diabetes patients with moderate control, excluding poorly controlled cases. | |
| KETO-1YR | Criteria: Diagnosis of ketoacidosis in the past year. |
| Explanation: Ketoacidosis is a serious condition with high blood ketones, often related to uncontrolled diabetes. Symptoms include nausea, confusion, and rapid breathing. | |
| MAJOR-DIABETES | Criteria: Major diabetes-related complications like amputation, kidney damage, or neuropathy. |
| Explanation: Includes significant health effects (e.g., nephropathy, retinopathy) stemming from uncontrolled diabetes. Multiple minor complications combined may also qualify. | |
| MAKES-DECISIONS | Criteria: Patient must make their own medical decisions. |
| Explanation: This assesses cognitive ability to understand medical options, communicate choices, and appreciate consequences. | |
| MI-6MOS | Criteria: Myocardial infarction (MI) within the past 6 months. |
| Explanation: MI refers to heart muscle damage from blocked blood flow. Assess via medical history and records confirming the event within 6 months. |